OpenStudy (anonymous):

What is the acceleration of a 50 kg grocery cart that is pushed with a net force of 12 N west

jimthompson5910 (jim_thompson5910):

f = m*a 12 = 50*a a = ??

OpenStudy (anonymous):

so 50 divided by 12?

OpenStudy (anonymous):

wait or do you have to figure out what times 50 equals 12??

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

but im not solving for force???

jimthompson5910 (jim_thompson5910):

to solve for 'a', you divide both sides by 50

OpenStudy (anonymous):

oh I got it now thanks

jimthompson5910 (jim_thompson5910):

you're welcome